# HUME: Measuring the Human-Model Performance Gap in Text Embedding Tasks

### v2

@@ D.2.2 Low-Agreement Tasks (Problematic Benchmarks)
Appendix E Additional Human vs Model Analysis
@@
Appendix F Models Evaluated
@@
Alibaba-NLP/gte-Qwen2-1.5B-instruct Li et al. (2023) 1780
google/embeddinggemma-300m Vera et al. (2025) 300
intfloat/e5-mistral-7b-instruct Wang et al. (2023; 2022) 7111
intfloat/multilingual-e5-large Wang et al. (2022) 560
intfloat/multilingual-e5-base Wang et al. (2022) 278
intfloat/multilingual-e5-small Wang et al. (2022) 118
mixedbread-ai/mxbai-embed-large-v1 Lee et al. (2024); Li & Li (2023) 335
NovaSearch/jasper_en_vision_language_v1 Zhang et al. (2025a) 1999
Qwen/Qwen3-Embedding-0.6B Zhang et al. (2025b) 596
Salesforce/SFR-Embedding-Mistral (Meng et al., 2024) 7110
sentence-transformers/all-MiniLM-L6-v2 Reimers & Gurevych (2019) 22.7
sentence-transformers/all-mpnet-base-v2 Reimers & Gurevych (2019) 109
@@
Appendix G LLM Usage Statement

### v3

@@ F.2.2 Low-Agreement Tasks (Problematic Benchmarks)
Appendix G Additional Human vs Model Analysis
@@
Appendix H Models Evaluated
@@
Alibaba-NLP/gte-Qwen2-1.5B-instruct li2023gte 1780
google/embeddinggemma-300m vera2025embeddinggemmapowerfullightweighttext 300
intfloat/e5-mistral-7b-instruct wang2023improving; wang2022e5 7111
intfloat/multilingual-e5-large wang2022e5 560
intfloat/multilingual-e5-base wang2022e5 278
intfloat/multilingual-e5-small wang2022e5 118
mixedbread-ai/mxbai-embed-large-v1 emb2024mxbai; li2023angle 335
NovaSearch/jasper_en_vision_language_v1 zhang2025jasperstelladistillationsota 1999
Qwen/Qwen3-Embedding-0.6B zhang2025qwen3embeddingadvancingtext 596
Salesforce/SFR-Embedding-Mistral (SFRAIResearch2024) 7110
sentence-transformers/all-MiniLM-L6-v2 reimers2019sentence 22.7
sentence-transformers/all-mpnet-base-v2 reimers2019sentence 109
@@
Appendix I Detailed LLM-as-Annotator Results
This section provides detailed LLM-as-annotator performance for each task-language pair. LLMs were not evaluated on clustering tasks due to the difficulty of eliciting cluster assignments from generative models.
Dataset Human GPT-5 GPT-4.1 Gemini Mistral Qwen3
Full Mini Full Mini 2.5 Flash Small-24B-I 30B 32B Coder
Emotion [eng] 45.8 37.5 33.3 41.7 29.2 35.4 43.1 37.5 38.3 42.1
Multilingual Sentiment [ara] 95.0 92.5 95.0 90.0 90.0 92.5 90.5 92.8 85.5 91.2
Multilingual Sentiment [eng] 77.5 100.0 100.0 97.5 95.0 97.5 97.2 99.0 93.0 96.5
Multilingual Sentiment [nob] 85.0 92.5 85.0 87.5 82.5 85.0 86.8 86.8 77.2 88.5
Multilingual Sentiment [rus] 92.5 87.5 85.0 75.0 85.0 92.5 64.0 74.2 75.8 78.0
Toxic Conversations [eng] 73.3 73.3 66.7 73.3 75.6 71.1 69.6 60.0 68.2 66.9
Tweet Sentiment [eng] 84.4 68.9 75.6 71.1 75.6 68.9 65.6 68.9 72.9 70.7
Average 79.1 78.9 77.2 76.6 76.1 77.6 73.8 74.2 73.0 76.3
Table 10: Detailed LLM-as-annotator results for classification tasks. Bold indicates best performance per row.
Dataset Human GPT-5 GPT-4.1 Gemini Mistral Qwen3
Full Mini Full Mini 2.5 Flash Small-24B-I 30B 32B Coder
Core17 [eng] 85.2 76.6 83.2 72.7 78.0 76.8 74.4 89.8 74.2 72.2
News21 [eng] 92.7 72.9 78.7 75.1 78.2 75.0 77.4 80.5 77.9 75.2
Robust04 [eng] 88.5 75.7 84.0 79.5 79.3 79.7 77.2 84.6 79.2 73.6
Wikipedia [dan] 91.4 84.7 78.3 85.8 83.6 82.8 84.0 76.6 79.8 79.9
Wikipedia [eng] 82.4 69.6 61.0 70.8 68.6 74.2 75.1 59.8 73.4 69.5
Wikipedia [nob] 89.8 71.5 68.1 70.6 75.8 68.6 79.6 62.6 64.2 72.3
Average 88.3 75.1 75.5 75.7 77.2 76.2 78.0 75.6 74.8 73.8
Table 11: Detailed LLM-as-annotator results for reranking tasks. Bold indicates best performance per row.
Dataset Human GPT-5 GPT-4.1 Gemini Mistral Qwen3
Full Mini Full Mini 2.5 Flash Small-24B-I 30B 32B Coder
SICK-R [eng] 82.6 68.5 67.5 66.9 72.9 59.0 66.7 59.4 57.3 68.8
STS12 [eng] 91.2 83.2 82.7 87.3 87.8 83.8 83.9 83.8 81.7 84.5
STS22 [ara] 67.5 57.1 45.5 56.9 55.8 43.1 56.9 35.3 50.4 47.9
STS22 [eng] 78.4 67.8 80.3 81.6 79.7 81.3 78.2 74.1 76.2 76.4
STS22 [rus] 58.7 71.9 51.3 67.2 66.6 64.9 78.6 65.9 62.4 67.0
STSBenchmark [eng] 80.4 89.3 86.7 89.7 86.3 83.8 85.4 84.3 84.0 82.9
Average 76.5 73.0 69.0 74.9 74.9 69.3 75.0 67.1 68.6 71.3
Table 12: Detailed LLM-as-annotator results for sts tasks. Bold indicates best performance per row.
I.1 Human-LLM Task Difficulty Correlation
To assess whether humans and LLMs face similar challenges, we computed Spearman rank correlations between human and LLM performance across the 19 task-language pairs (excluding clustering). A positive correlation indicates that tasks where humans perform well also tend to be tasks where LLMs perform well.
LLM Model Spearman $\rho$ $p$-value
GPT-5 Full 0.47 0.042
GPT-5 Mini 0.52 0.023
GPT-4.1 Full 0.43 0.066
GPT-4.1 Mini 0.57 0.012
Gemini 2.5 Flash 0.50 0.029
Mistral Small 0.30 0.215
Qwen3-30B 0.50 0.031
Qwen3-32B 0.52 0.022
Qwen3-Coder 0.57 0.011
Average (all LLMs) 0.52 0.023
Table 13: Spearman rank correlation between human and LLM performance across 19 task-language pairs (clustering excluded). A positive $\rho$ indicates that tasks where humans score high also tend to be tasks where LLMs score high. Values with $p<0.05$ indicate statistically significant correlations. The moderate positive correlations suggest partially shared task difficulty patterns between humans and LLMs.
The overall correlation is moderate and statistically significant ($\rho=0.52$, $p=0.023$), suggesting partially shared difficulty patterns.
Appendix J LLM Usage Statement